# Semiclassical Mixmaster Universe

### v1

@@ I Introduction
The behavior of the universe close to the big bang singularity has been a question of much interest over the past decades. Despite several notable works, a widely accepted picture has yet to emerge. Seminal papers by Belinski, Khalitnikov, and Lifshitz, Belinskii and Khalatnikov (1969); Belinskiĭ et al. (1971) showed that the approach to a spacelike singularity in Einstein gravity is homogeneous, dominated by time derivatives, and therefore described as a dynamical system of coupled non-linear ordinary differential equations (ODE). This led to the study of homogeneous anisotropic Bianchi models that describe the evolution of scale factors moving in potentials that come from spatial curvature Ryan and Shepley (1975). The most interesting of these is the Bianchi IX model, also termed the “Mixmaster Universe” Misner (1969a). It describes the universe as a billiard ball in a triangular potential with exponentially high expanding wall. Between the walls, the potential is negligible and the universe is described by the Kasner spacetime
$ds^{2}=-N^{2}dt^{2}+\sum_{i=1}^{3}t^{2p_{i}}dx_{i}^{2},$ (1)
where the so-called Kasner exponents $p_{i}$ satisfy the sum rules $\sum_{i}p_{i}=1$ and $\sum_{i}p_{i}^{2}=1$ by virtue of the Einstein equations. These exponents transition to new ones at each collision with a wall of the potential, sending the universe into a new Kasner epoch with different exponents. The transitions are described by rules that have been the subject of many studies in vacuum Misner (1969b) and with matter coupling Ali and Husain (2017a). This interesting dynamics exhibits chaos, a feature that has led to many investigations over the years Chernoff and Barrow (1983); Barrow (1981); Hobill et al. (1991); Burd et al. (1990); Kiefer et al. (2018); Bini et al. (2009); Andriopoulos and Leach (2008).
@@ III.1 Volume, anisotropy and Kasner sum evolution
Anisotropies: The second row compares the anisotropies $\beta_{\pm}$ in each of the cases. The approach to the first bounce is noteworthy: all cases are very similar as the first bounce is approached, with anisotropies vanishing near $t=40$. However after the bounce, the second and third frames (the cases $H^{\rm{eff}}_{C}$ and $H^{\rm{eff}}_{Q}$) show higher anisotropies that rise rapidly near each bounce; it is also evident that the classical and quantum scalar field cases differ significantly in this respect. Fig. 2 is a view of the anisotropies at the first bounce point of Fig. 1, showing that all three bounce cases are qualitatively similar.
@@
III.2 Scalar field excitation
@@
Lastly Fig. 4 shows a zoomed in version of the even state excitations of the scalar field for $H_{q}^{\rm{eff}}$ near the first bounce; the left frame shows the lower levels, and the right frame shows the higher levels near the truncation $n=30$; the log of the volume (dashed red line) is overlaid. This shows order $10^{-2}$ excitations of the higher levels with a similar pattern.
@@ IV Summary and Discussion
The observation that chaos is reduced in effective gravity is not surprising, since intuitively, chaos is produced by the increasing number of bounces off the triangular potential walls as the singularity is approached. If the universe achieves a minimum volume, the potential walls stop closing in, and it is reasonable to expect that the number of collisions with the walls reaches a maximum, and then starts to reduce through the bounce. The presence of matter affects this only in the transfer of some “energy” between the matter and gravity terms in the Hamiltonian constraint; the reduction in chaos appears to be mainly due to the bounce in the effective gravity cases. This argument also suggests that chaos is not merely a coordinate artefact, and that although Lyapunov exponents may differ if other phase space variables and time coordinates are used, it appears unlikely that all exponents would become zero or negative, given the nature of the potential.
This work extends the much-studied vacuum Bianchi IX model in two ways – by introducing effective gravity from fundamental discreteness and quantum matter with back reaction. The same methods may be used for the other Bianchi models, with their somewhat simpler potentials. But more interesting from a quantum gravity perspective are inhomogeneous models, the simplest of which are the Gowdy cosmologies with one local degree of freedom. Introducing quantum matter with back reaction in the manner we have done here would require introducing matter expectation value terms in both the hamiltonian and diffeomorphism constraints, and ensuring their algebra remains consistent. This remains to be seen.

### v2

@@ I Introduction
The behavior of the universe close to the big bang singularity has been a question of much interest over the past decades. Despite several notable works, a widely accepted picture has yet to emerge. Seminal papers by Belinski, Khalitnikov, and Lifshitz, Belinskii and Khalatnikov (1969); Belinskiĭ et al. (1971) showed that the approach to a spacelike singularity in Einstein gravity is homogeneous, dominated by time derivatives, and therefore described as a dynamical system of coupled non-linear ordinary differential equations (ODE). This led to the study of homogeneous anisotropic Bianchi models that describe the evolution of scale factors moving in potentials that come from spatial curvature Ryan and Shepley (1975). The most interesting of these is the Bianchi IX model, also termed the “Mixmaster Universe” Misner (1969a). It describes the universe as a billiard ball in a triangular potential with exponentially high expanding wall. Between the walls, the potential is negligible and the universe is described by the Kasner spacetime $ds^{2}=-N^{2}dt^{2}+\sum_{i=1}^{3}t^{2p_{i}}dx_{i}^{2}$, where the so-called Kasner exponents $p_{i}$ satisfy the sum rules $\sum_{i}p_{i}=1$ and $\sum_{i}p_{i}^{2}=1$ by virtue of the Einstein equations. These exponents transition to new ones at each collision with a wall of the potential, sending the universe into a new Kasner epoch with different exponents. The transitions are described by rules that have been the subject of many studies in vacuum Misner (1969b) and with matter coupling Ali and Husain (2017a). This interesting dynamics exhibits chaos, a feature that has led to many investigations over the years Chernoff and Barrow (1983); Barrow (1981); Hobill et al. (1991); Burd et al. (1990); Kiefer et al. (2018); Bini et al. (2009); Andriopoulos and Leach (2008).
@@ III.1 Volume, anisotropy and Kasner sum evolution
Anisotropies: The second row compares the anisotropies $\beta_{\pm}$ in each of the cases. The approach to the first bounce is noteworthy: all cases are very similar as the first bounce is approached, with anisotropies vanishing near $t=40$. However after the bounce, the second and third frames (the cases $H^{\rm{eff}}_{C}$ and $H^{\rm{eff}}_{Q}$) show higher anisotropies that rise rapidly near each bounce; it is also evident that the classical and quantum scalar field cases differ significantly in this respect. Fig. 2 is a view of the anisotropies at the first bounce point of Fig. 1; it shows that all three bounce cases are qualitatively similar. The evolutions of $\beta_{+}$ and $\beta_{-}$ are quite different due to their definitions in (3) and (4): the former is a function of all three scale factors, while the latter depends on only two. (The different evolutions are a feature also of the exact Kasner solution where $\beta_{+}\sim\ln(t^{p_{1}+p_{2}-p_{3}^{2}})$ and $\beta_{+}\sim\ln(t^{p_{1}-p_{2}})$, for Kasner exponents $p_{1},p_{2},p_{3}$ satisfying the sum rule.)
@@
However, the orange curves (corresponding to $H_{IX}^{\rm{eff}},\ H_{C}^{\rm{eff}}$ and $H_{Q}^{\rm{eff}}$ —where there is a bounce) show that the sum rule is a constant away from the bounce point, but not unity due to the polymer scale $\lambda=0.1$; it undergoes oscillations through the bounce (where the green volume curve dips); the differences in the three frames in this row show that “matter matters” when the universe bounces, and in particular whether matter is classical or quantum. Furthermore, the oscillations in the sum rule indicate that the approach to the volume bounce is not Kasner. This is in contrast with the BKL conjecture, which states that the approach to the classical singularity is Kasner-like; the conjecture is of course silent on the singularity free bouncing cosmologies we study here.
III.2 Scalar field excitation and density
@@
Fig. 4 shows a zoomed in version of the even state excitations of the scalar field for $H_{q}^{\rm{eff}}$ near the first bounce; the left frame shows the lower levels, and the right frame shows the higher levels near the truncation $n=30$; the log of the volume (dashed red line) is overlaid. This shows order $10^{-2}$ excitations of the higher levels with a similar pattern.
@@
Lastly, Fig. 6 exhibits the conservation of the Hamiltonian constraint for each of the cases we study to order at least to order $10^{-7}$. This shows the accuracy of our numerical evolutions, including the oscillations of the Kasner sum at the bounce point in Fig. 1. It also reflects probability conservation of the quantum scalar field case (third frame), where any probability violation would result in non-conservation of the Hamiltonian constraint through the expectation value of the matter contribution to this constraint.
@@ III.3.2 Fractal dimension
Our main result of this subsection is evidence of significant chaos reduction in the cases where there is a bounce, as exhibited in Table 1. These results are intuitively reasonable from the perspective that a minimum volume effectively saturates the number of collisions of the universe particle with the walls of the Bianchi IX potential, compared to the case where there is no minimal volume and the singularity is present. This in turn means that a more pervasive exploration of the phase space is thwarted by the bounce in volume, thereby reducing the degree of chaos. As this intuition is broad, it likely applies to a͡ny mechanism that replaces the singularity with a bounce; see e.g. Moriconi et al. (2014); Barca and Giovannetti (2025) for other approaches where chaos is either reduced or eliminated altogether.
@@ IV Summary and Discussion
As noted above, the observation that chaos is reduced in effective gravity is not surprising, since chaos is produced by the increasing number of bounces off the triangular potential walls as the singularity is approached. If the universe achieves a minimum volume, the potential walls stop closing in, and it is reasonable to expect that the number of collisions with the walls reaches a maximum, and then starts to reduce through the bounce. The presence of matter affects this only in the transfer of some “energy” between the matter and gravity terms in the Hamiltonian constraint; the reduction in chaos appears to be mainly due to the bounce in the effective gravity cases. This argument also suggests that chaos is not merely a coordinate artefact, and that although Lyapunov exponents may differ if other phase space variables and time coordinates are used, it appears unlikely that all exponents would become zero or negative, given the nature of the potential.
This work extends the much-studied vacuum Bianchi IX model in two ways – by introducing effective gravity via fundamental discreteness corrections and quantum matter with back reaction. The same methods may be used for the other Bianchi models, with their somewhat simpler potentials. More interesting from a quantum gravity perspective are inhomogeneous models, the simplest of which are the Gowdy cosmologies with one local degree of freedom. Introducing quantum matter with back reaction in the manner we have done here would require introducing matter expectation value terms in both the hamiltonian and diffeomorphism constraints, and ensuring their algebra remains consistent. This remains to be seen.
Acknowledgements This work was supported by the Natural Science and Engineering Research Council of Canada.